# Joint Beamforming and Resource Allocation for Delay Optimization in RIS-Assisted OFDM Systems: A DRL Approach

### v1

@@ Abstract
This paper investigates a joint phase design and resource allocation problem in downlink reconfigurable intelligent surface (RIS)-assisted orthogonal frequency division multiplexing (OFDM) systems to optimize average delay, where data packets for each user arrive at the base station stochastically. The sequential optimization problem is inherently a Markov decision process (MDP), making it fall within the scope of reinforcement learning. To effectively handle the mixed action space and reduce the state space dimensionality, a hybrid deep reinforcement learning (DRL) approach is proposed. Specifically, proximal policy optimization (PPO)-$\Theta$ is employed to optimize RIS phase shift design, while PPO-N is responsible for subcarrier allocation decisions. To further mitigate the curse of dimensionality associated with subcarrier allocation, a multi-agent strategy is introduced to optimize subcarrier allocation indicater more efficiently. Moreover, to achieve more adaptive resource allocation and accurately capture network dynamics, key factors closely related to average delay, including the number of backlogged packets in buffers and the current packet arrivals, are incorporated into the state space. Furthermore, a transfer learning framework is introduced to enhance training efficiency and accelerate convergence. Simulation results demonstrate that the proposed algorithm significantly reduces average delay, enhances resource allocation efficiency, and achieves superior system robustness and fairness compared to baseline methods.
@@ I Introduction
However, given the inherently limited nature of wireless resources, ensuring low-latency communication in RIS-assisted wireless systems remains a fundamental challenge, where the phase design and resource allocation are coupled. In addition to the challenge caused by joint optimization of phase design and resource allocation, limited spectrum and power resources constrain the system’s ability to support high data rates and real-time responsiveness, multi-path effects introduce rapid fluctuations in channel conditions and signal quality [6], which necessiates network management to be adaptive to the environment. On the other hand, different users may have different numbers of backlogged packets, which needs to be taken into account in delay-aware resource allocation [7]. Therefore, in delay-aware RIS-assisted wireless networks, it is critical to make sequential resource allocation and phase design decisions while being adaptive to real-time channel conditions and queue status [8], which is challenging due to complex network environment and mixed decision variables including continuous power and discrete subcarrier. To this end, the intelligent and adaptive management of communication resources, capable of accommodating dynamic network conditions and fluctuating traffic demands, has become a critical research focus for achieving ultra-low latency and high reliability in future wireless networks.
@@ I-A Prior Works and Motivation
To address delay-aware resource allocation in wireless networks, some studies have utilized large-scale fading information for spectrum and power allocation. For example, in device-to-device (D2D)-enabled vehicular networks, resource allocation schemes based on large-scale fading were proposed to enhance the ergodic capacity of vehicle-to-infrastructure (V2I) links while ensuring vehicle-to-vehicle (V2V) communication reliability [12]. Similarly, a joint power and rate allocation strategy was developed for single-input multiple-output (SIMO) systems under ultra-reliability and energy efficiency constraints, relying only on average signal and interference statistics [13]. In addition, Sun et al. designed a separate resource block allocation and power control algorithm for D2D-based V2V communications, effectively meeting latency and reliability constraints without requiring instantaneous channel state information (CSI) [14]. However, these static approaches fail to capture fast channel variations, limiting their capacity. To overcome this, researchers have increasingly adopted the Lyapunov optimization framework, which enables real-time delay-aware resource allocation by jointly considering queue dynamics, channel fluctuations, and system constraints. For instance, Guo et al. proposed a Lyapunov-based cross-layer design integrating transport-layer rate control with physical-layer allocation to meet delay constraints and mitigate cross-tier interference [15]. Similar methods have been applied to develop robust scheduling policies that ensure queue stability and service quality in dynamic environments [16, 17, 18], marking a shift from static optimization to adaptive, queue-aware strategies for low-latency wireless communication.
@@
While RIS has shown great promise in enhancing spectral efficiency, coverage, and energy efficiency [22, 23, 24, 25, 26], its application to latency-sensitive scenarios remains limited. Existing studies primarily focus on mobile edge computing (MEC) settings and often overlook queuing delay [27, 28, 29], which is critical for real-time communications. For instance, [27] proposed an iterative algorithm to minimize task delay and energy consumption in an RIS-assisted non-orthogonal multiple access (NOMA)-MEC network, while [28] considered joint delay and rate optimization in RIS-assisted multiple-input multiple-output (MIMO)-MEC systems. To address these limitations, recent studies have applied Lyapunov-based techniques for queue-aware resource allocation in RIS-assisted systems. By transforming long-term delay and stability objectives into tractable per-slot problems, these methods enable dynamic adaptation to traffic fluctuations [30, 31]. For example, in [31], an RIS-assisted semi-grant-free NOMA system was optimized through joint design of power, channel assignment, and reflection coefficients, achieving improved queue stability over RIS-free baselines in RIS-assisted wireless communication systems. Nevertheless, delay-aware resource allocation in RIS-assisted wireless systems is still far from being fully addressed. Existing Lyapunov-based methods often rely on complex alternating optimization procedures and typically do not explicitly incorporate queuing delay into the objective, limiting their practical applicability in delay-sensitive scenarios.
@@ I-B Contributions and Organization
This paper aims to investigate resource allocation methods for delay optimization in RIS-assisted networks by leveraging DRL, with a particular focus on maintaining user fairness. To achieve this goal, a hybrid DRL framework is proposed to tackle the challenge of delay optimization in such networks. The state space is carefully designed by incorporating various factors related to communication quality to achieve excellent delay performance while maintaining fairness. Meanwhile, transfer learning is employed to accelerate the convergence of the learning process. The main contributions of this work are summarized as follows.
  • We adopt a hybrid DRL method, where the central controller at the BS acts as an agent and is trained using the proximal policy optimization (PPO) algorithm. Specifically, a continuous-action PPO-$\Theta$ is used to determine the RIS phase shift design, while a discrete-action PPO-N is employed for subcarrier allocation decisions. Moreover, to alleviate the curse of dimensionality caused by the multi-dimensional discrete variables of subcarrier allocation indicators, we adopt a discrete-action multi-agent approach to optimize subcarrier allocation, where each agent is responsible for the allocation of a single subcarrier.
  • To enhance delay performance while maintaining fairness and ensuring the robustness of the algorithm in adapting to diverse scenarios, the state is defined to include channel state information, the number of currently arriving data packets and the number of backlogged packets in buffers. This information helps capture both the channel conditions and the users’ delay status, enabling more effective resource allocation in dynamic RIS-assisted networks.
@@ II System Model and Problem formulation
We consider an RIS-assisted downlink multiple-input single-output (MISO) orthogonal frequency division multiplexing (OFDM) communication system, as illustrated in Fig. 1LABEL:sub@Fig:RIS, where a base station (BS) equipped with $N_{t}$ antennas serves $K$ single-antenna users. An RIS comprising $M$ passive reflecting elements is deployed to effectively enhance the transmission quality. The entire communication duration is divided into $T_{\mathrm{sum}}$ timeslots, indexed by $\{1,2,\ldots,t,\ldots,T_{\mathrm{sum}}\}$, each with a fixed length of $T$. Within each time slot, the wireless channel is assumed to remain approximately static. Each user’s data packets, which arrive randomly at the corresponding buffer at the BS, follow a first-come, first-served (FCFS) discipline, as illustrated in Fig. 1LABEL:sub@Fig:queue. In each time slot, the accumulated packets in the buffer $k$ are scheduled for transmission to the user $k$. Packets that can not be delivered will remain in the buffer and continue to accumulate, resulting in increased packet delay over time.
@@ II-A Signal Model
where $\tilde{\mathbf{G}}_{l}[t]=\mathbf{0}\text{,}\;l\in\begin{Bmatrix}1-L_{2}$\text{\,} $\ldots,-1\end{Bmatrix}\cup\begin{Bmatrix}L_{1},\ldots,L-1\end{Bmatrix}$, and $\boldsymbol{\Phi}[t]=\operatorname{diag}(e^{j\theta_{1}[t]},e^{j\theta_{2}[t]},...,e^{j\theta_{M}[t]})\in\mathbb{C}^{M\times M}$ is the RIS reflection coefficient matrix at the time slot $t$. Let $\boldsymbol{\mathbf{\varphi}}[t]=[e^{j\theta_{1}[t]},e^{j\theta_{2}[t]},...,e^{j\theta_{M}[t]}]^{T}\in\mathbb{C}^{M\times 1}$ denote the RIS reflection coefficient vector at the time slot $t$, where $\theta_{m}[t]$ refers to the phase shift of the reflection element $m$ at the time slot $t$. We set $\tilde{\mathbf{h}}^{(r)}_{l,k}[t]=\sum_{i=0}^{L_{2}-1}\mathrm{diag}(\tilde{\mathbf{r}}_{i,k}[t])\tilde{\mathbf{G}}_{l-i}[t]$, thus $\tilde{\mathbf{h}}_{l,k}[t]$ can further be expressed as
@@ II-C Problem Formulation
In this paper, our objective is to minimize the average delay by jointly optimizing the subcarrier allocation shift $\alpha_{n,k}[t]$, beamforming vectors at the BS $\mathbf{w}_{n}[t]$, and RIS reflection coefficient $\theta_{m}[t]$. Specifically, the problem is formulated as
@@ III Deep Reinforcement Learning Based Algorithm
This section introduces a DRL-based framework for addressing the problem in (11), highlighting the core design principles and the corresponding training methodology.
@@ III-A Overall Framework
The optimization problem presented in equation (11) necessitates determining the RIS phase shifts and resource allocation for each time slot, where decisions made in the current slot directly influence the number of packets that can be delivered, thereby affecting the queue length in the subsequent slot. Therefore, the problem (11) is a sequential decision making issue, which can be typically characterized by MDP. To this end, a hybrid RL approach is adopted, wherein a central controller collects information regarding the channel state information, users’ rates, and the states of all buffers. The central controller, deployed at the BS, functions as an agent responsible for controlling the RIS and subcarrier allocation. It explores the unknown communication environment to acquire experience, which is subsequently utilized to guide policy design and optimize the RIS reflection phase shift and subcarrier allocation strategies based on observations of the environment state. Specifically, the PPO-$\Theta$ is utilized to predict the RIS reflection phase shift, while PPO-$N$ is applied to determine the subcarrier allocation indicaters. The detailed algorithm framework is illustrated in Fig. 2.
@@ III-B PPO-$\Theta$ Design
The reward function plays a pivotal role in DRL, as it directly steers the agent’s learning behavior toward the desired optimization objective. To effectively minimize the average delay in the system, it is essential to align the reward function with the optimization goal.
To this end, the proposed framework defines the reward at each time step as the negative total number of packets queued in all users’ buffers, formulated as
@@
(c) Critic network architecture of PPO-N
@@ III-C PPO-$N$ Design
The PPO-$N$ framework is employed to optimize subcarrier allocation indicators. Since subcarrier allocation indicators are discrete variables, a discrete PPO algorithm is utilized for optimization. However, if a single-agent approach is adopted to jointly output all subcarrier allocation indicators, the action space becomes prohibitively large. Specifically, allocating $N$ subcarriers to $K$ users results in $K^{N}$ possible allocation combinations, leading to an action space dimension of $K^{N}$ for a single agent, which causes the curse of dimensionality. To mitigate this issue, a multi-agent approach is adopted to decompose the action space. Specifically, each agent is responsible for determining the allocation indicator of a specific subcarrier. The system thus comprises $N$ agents, each with an action space dimension of $K$, significantly reducing the complexity. Moreover, this design inherently satisfies constraints (11a) and (11b).
@@
The global state $s_{t}^{(N)}$ used by the centralized critic includes the complete equivalent channel state information $\mathbf{h}_{n,k}^{\text{eff}}$ for all users and subcarriers, along with the global buffer state $q_{k}[t]$ and packet arrivals $\ell_{k}[t]$, enabling joint policy evaluation across agents. To reduce input dimensionality, $\mathbf{h}_{n,k}^{\text{eff}}$ is defined as $\mathbf{\bar{h}}^{(d)}_{n,k}+\mathbf{\varphi}^{T}\mathbf{\bar{h}}^{(r)}_{n,k}$, where $\mathbf{\varphi}[t]$ is the RIS reflection coefficient vector generated by PPO-$\Theta$. The local observation $o_{t}^{(n)}$ of agent $n$ comprises the equivalent channel state from the BS to users on subcarrier $n$, buffer length $q_{k}[t]$, and packet arrivals $\ell_{k}[t]$.
Since the objective is to minimize average delay by jointly optimizing RIS phase shifts and subcarrier allocation, PPO-$N$ and PPO-$\Theta$ share the reward function defined in (12). As $N$ agents allocate $N$ subcarriers with $K$ possible user choices, the action space $a_{t}^{(n)}$ of each agent is $\{1,2,\dots,K\}$.
In this algorithm, since $N$ agents are used to allocate $N$ subcarriers, and each subcarrier has $K$ possible allocation options, the action space $a_{t}^{(n)}$ of each agent is defined as $\{1,2,\dots,K\}$.
@@
b) Critic network design: The input to the centralized critic network corresponds to the global state space $s_{t}^{(N)}$. This input is fed into a fully connected neural network to estimate the value of the global state. The critic network outputs a scalar value $V_{\phi^{(N)}}(s_{t}^{(N)})$, which approximates the expected return under the current policy. The detailed architecture of the critic network is illustrated in Fig. 3LABEL:sub@PPONcritic.
@@ III-F Computational Complexity
Figure 4: Illustration of the simulated scenario.
IV Numerical Results
IV-A Simulation Settings
[…]
@@ IV-C Convergence and Validity Analysis
Figure 7: Impact of reflection elements $M$ on delay performance.
Figure 8: Average delay performance as arrival rate $\lambda$ varies in brustiness scenarios.

### v3

# Beamforming and Resource Allocation for Delay Optimization in RIS-Assisted OFDM Systems

@@ Abstract
This paper investigates a joint beamforming and resource allocation problem in downlink reconfigurable intelligent surface (RIS)-assisted orthogonal frequency division multiplexing (OFDM) systems to optimize average delay, where data packets for each user arrive at the base station stochastically. The sequential optimization problem is inherently a Markov decision process (MDP), thus falling within the scope of reinforcement learning. To effectively handle the mixed action space and reduce the state space dimensionality, a hybrid deep reinforcement learning (DRL) approach is proposed. Specifically, proximal policy optimization (PPO)-$\Theta$ is employed to optimize the RIS phase shift design, while PPO-$N$ is responsible for subcarrier allocation decisions. The active beamforming at the base station is then derived from the jointly optimized RIS phase shifts and subcarrier allocation decisions. To further mitigate the curse of dimensionality associated with subcarrier allocation, a multi-agent strategy is introduced to optimize subcarrier allocation indicators more efficiently. Moreover, to achieve more adaptive resource allocation and accurately capture network dynamics, key factors closely related to average delay—such as the number of backlogged packets in buffers and current packet arrivals—are incorporated into the state space. Furthermore, a transfer learning framework is introduced to enhance training efficiency and accelerate convergence. Simulation results demonstrate that the proposed algorithm significantly reduces average delay, enhances resource allocation efficiency, and achieves superior system robustness and fairness compared to baseline methods.
@@ I Introduction
However, given the inherently limited nature of wireless resources, ensuring low-latency communication in RIS-assisted wireless systems remains a fundamental challenge, where the phase design and resource allocation are coupled. In addition to the challenge caused by joint optimization of phase design and resource allocation, limited spectrum and power resources constrain the system’s ability to support high data rates and real-time responsiveness, multi-path effects introduce rapid fluctuations in channel conditions and signal quality [6], which necessiates network management to be adaptive to the environment. On the other hand, different users may have different numbers of backlogged packets, which needs to be taken into account in delay-aware resource allocation [7]. Therefore, in delay-aware RIS-assisted wireless networks, it is critical to make sequential resource allocation and phase design decisions while being adaptive to real-time channel conditions and queue status [8], which is challenging due to complex network environment and mixed decision variables including continuous RIS phase shifts design and discrete subcarrier.
@@ I-A Prior Works and Motivation
To address delay-aware resource allocation, some works leverage large-scale fading for spectrum and power management. For instance, in device-to-device (D2D)-enabled vehicular networks, resource allocation based on large-scale fading improves vehicle-to-infrastructure (V2I) ergodic capacity while ensuring vehicle-to-vehicle (V2V) reliability [12]. Similarly, single-input multiple-output (SIMO) systems have adopted power and rate allocation strategies using average signal and interference statistics under reliability and energy efficiency constraints [13]. Sun et al. further proposed a resource block and power control algorithm for D2D-based V2V communications without relying on instantaneous channel state information (CSI) [14]. However, these static methods struggle to adapt to rapid channel variations. To address this, Lyapunov optimization has gained traction for real-time delay-aware resource management, accounting for queue dynamics, channel fluctuations, and system constraints. For example, Guo et al. developed a cross-layer Lyapunov framework combining transport-layer rate control and physical-layer allocation to meet delay targets [15]. Similar approaches have led to robust scheduling policies that ensure queue stability and service quality in dynamic environments [16, 17, 18].
@@
While RIS has enhanced spectral efficiency, coverage, and energy efficiency [22, 23, 24, 25, 26], its application in latency-sensitive scenarios remains limited. Most existing works focus on mobile edge computing and neglect queuing delay, which is crucial for real-time communication [27, 28, 29]. For instance, [27] minimized task delay and energy in RIS-assisted non-orthogonal multiple access networks, and [28] addressed joint delay and rate in multiple-input multiple-output mobile edge computing systems. To address this gap, recent studies adopt Lyapunov-based methods for queue-aware resource allocation, converting long-term delay goals into per-slot decisions to adapt to dynamic traffic [30, 31]. For example, [31] jointly optimized power, channel assignment, and reflection coefficients in a semi-grant-free RIS-assisted system, improving queue stability over RIS-free baselines. However, current approaches remain insufficient. Many rely on complex alternating optimization and fail to explicitly include queuing delay in the objective, limiting their effectiveness in truly delay-sensitive RIS-assisted systems.
@@ I-B Contributions and Organization
This paper aims to investigate resource allocation methods for delay optimization in RIS-assisted networks by leveraging DRL, with a particular focus on maintaining user fairness. To achieve this goal, a hybrid DRL framework is proposed to tackle the challenge of delay optimization in such networks. The main contributions of this work are summarized as follows.
  • A hybrid DRL framework is adopted, where the base station (BS) controller acts as the agent and is trained using proximal policy optimization (PPO). Specifically, PPO-$\Theta$ with continuous action space is used for RIS phase shift optimization, while PPO-$N$ with discrete action space handles subcarrier allocation. To address the curse of dimensionality, a multi-agent strategy is employed, where each agent is responsible for the allocation of a single subcarrier.
  • To ensure low delay, fairness, and adaptability, the state space incorporates CSI, the number of arriving packets, and the number of backlogged packets in each buffer. This enables the agent to capture both channel conditions and delay status for more efficient resource allocation in dynamic environments.
@@ II System Model and Problem formulation
We consider an RIS-assisted downlink multiple-input single-output (MISO) orthogonal frequency division multiplexing (OFDM) communication system, as illustrated in Fig. 1LABEL:sub@Fig:RIS, where a BS equipped with $N_{t}$ antennas serves $K$ single-antenna users. An RIS comprising $M$ passive reflecting elements is deployed to effectively enhance the transmission quality. The entire communication duration is divided into $T_{\mathrm{sum}}$ timeslots, indexed by $\{1,2,\ldots,t,\ldots,T_{\mathrm{sum}}\}$, each with a fixed length of $T$. Within each time slot, the wireless channel is assumed to remain approximately static. Each user’s data packets, which arrive randomly at the corresponding buffer at the BS, follow a first-come, first-served (FCFS) discipline, as illustrated in Fig. 1LABEL:sub@Fig:queue. In each time slot, the accumulated packets in the buffer $k$ are scheduled for transmission to the user $k$. Packets that can not be delivered will remain in the buffer and continue to accumulate, resulting in increased packet delay over time.
@@ II-A Signal Model
where $\tilde{\mathbf{G}}_{l}[t]=\mathbf{0}\text{,}\;l\in\begin{Bmatrix}1-L_{2}$\text{\,} $\ldots,-1\end{Bmatrix}\cup\begin{Bmatrix}L_{1},\ldots,L-1\end{Bmatrix}$, and $\boldsymbol{\Phi}[t]=\operatorname{diag}(e^{j\theta_{1}[t]},e^{j\theta_{2}[t]},...,e^{j\theta_{M}[t]})\in\mathbb{C}^{M\times M}$ is the RIS reflection phase shifts matrix at the time slot $t$. Let $\boldsymbol{\mathbf{\varphi}}[t]=[e^{j\theta_{1}[t]},e^{j\theta_{2}[t]},...,e^{j\theta_{M}[t]}]^{T}\in\mathbb{C}^{M\times 1}$ denote the RIS reflection phase shifts vector at the time slot $t$, where $\theta_{m}[t]$ refers to the phase shift of the reflection element $m$ at the time slot $t$. We set $\tilde{\mathbf{h}}^{(r)}_{l,k}[t]=\sum_{i=0}^{L_{2}-1}\mathrm{diag}(\tilde{\mathbf{r}}_{i,k}[t])\tilde{\mathbf{G}}_{l-i}[t]$, thus $\tilde{\mathbf{h}}_{l,k}[t]$ can further be expressed as
@@ II-C Problem Formulation
In this paper, our objective is to minimize the average delay by jointly optimizing the subcarrier allocation shift $\alpha_{n,k}[t]$, beamforming vectors at the BS $\mathbf{w}_{n}[t]$, and RIS reflection phase shifts $\theta_{m}[t]$. Specifically, the problem is formulated as
@@ III Deep Reinforcement Learning Based Algorithm
This section introduces a DRL-based framework for addressing the problem in (11), highlighting the core design principles, the corresponding training methodology and the computational complexity of proposed algorithm.
@@ III-A Overall Framework
The optimization problem presented in equation (11) necessitates determining the RIS phase shifts and resource allocation for each time slot, where decisions made in the current slot directly influence the number of packets that can be delivered, thereby affecting the queue length in the subsequent slot. Therefore, the problem (11) is a sequential decision making issue, which can be typically characterized by MDP. To this end, a hybrid RL approach is adopted, wherein a central controller collects information regarding the channel state information and the states of all buffers. The central controller, deployed at the BS, functions as an agent responsible for controlling the RIS and subcarrier allocation. It explores the unknown communication environment to acquire experience, which is subsequently utilized to guide policy design and optimize the RIS reflection phase shift and subcarrier allocation strategies based on observations of the environment state. Specifically, the PPO-$\Theta$ is utilized to predict the RIS reflection phase shift, while PPO-$N$ is applied to determine the subcarrier allocation indicaters. The detailed algorithm framework is illustrated in Fig. 2.
@@ III-B PPO-$\Theta$ Design
The reward function plays a pivotal role in DRL, as it directly steers the agent’s learning behavior toward the desired optimization objective. To effectively minimize the average delay in the system, it is essential to align the reward function with the optimization goal. To this end, the proposed framework defines the reward at each time step as the negative total number of packets queued in all users’ buffers, formulated as
@@
(c) Centralized critic network architecture of PPO-$N$
@@ III-C PPO-$N$ Design
The PPO-$N$ framework optimizes discrete subcarrier allocation indicators using a discrete PPO algorithm. A single-agent approach faces a huge action space of size $K^{N}$ when allocating $N$ subcarriers to $K$ users, causing the curse of dimensionality. To address this, a multi-agent approach is used, where each of the $N$ agents controls one subcarrier’s allocation with an action space of size $K$, greatly reducing complexity. This design also naturally satisfies constraints (11a) and (11b).
@@
The global state $s_{t}^{(N)}$ used by the centralized critic includes the complete equivalent channel state information $\mathbf{h}_{n,k}^{\text{eff}}$ for all users and subcarriers, along with the global buffer state $q_{k}[t]$ and packet arrivals $\ell_{k}[t]$, enabling joint policy evaluation across agents. To reduce input dimensionality, $\mathbf{h}_{n,k}^{\text{eff}}$ is defined as $\mathbf{\bar{h}}^{(d)}_{n,k}+\mathbf{\varphi}^{T}\mathbf{\bar{h}}^{(r)}_{n,k}$, where $\mathbf{\varphi}[t]$ is the RIS reflection phase shifts vector generated by PPO-$\Theta$. The local observation $o_{t}^{(n)}$ of agent $n$ comprises the equivalent channel state from the BS to users on subcarrier $n$, buffer length $q_{k}[t]$, and packet arrivals $\ell_{k}[t]$.
Since the objective is to minimize average delay by jointly optimizing RIS phase shifts and subcarrier allocation, PPO-$N$ and PPO-$\Theta$ share the reward function defined in (12). Since $N$ agents are used to allocate $N$ subcarriers, and each subcarrier has $K$ possible allocation options, the action space $a_{t}^{(n)}$ of each agent is defined as $\{1,2,\dots,K\}$.
@@
b) Centralized critic network design: The input to the centralized critic network corresponds to the global state space $s_{t}^{(N)}$. This input is fed into a fully connected neural network to estimate the value of the global state. The critic network outputs a scalar value $V_{\phi^{(N)}}(s_{t}^{(N)})$, which approximates the expected return under the current policy. The detailed architecture of the critic network is illustrated in Fig. 3LABEL:sub@PPONcritic.
@@ IV-B Baseline Methods
Figure 7: Average delay performance as packet arrival rate $\lambda$ varies.